# DiverseDream: Diverse Text-to-3D Synthesis with Augmented Text Embedding

### v1

@@ 5.1 Comparison with SOTA
Qualitative comparison. The comparative results in Fig. 4 demonstrate that our method offers more diversity among particles compared to VSD and SDS. For example, when given “A high-quality ice cream sundae” prompts, VSD tends to collapse into cone-shaped ice cream, while our method is capable of generating glass shapes and other variants.
Furthermore, we observe that our generated 3D models inherit the texture and structure of the reference image, as depicted in Fig. 1. This indicates the potential of transferring the diversity of 2D to 3D through text prompt augmentation. It is worth noting that we do not require the HiPer framework to perfectly invert the reference images, we only need HiPer tokens to capture the essence of reference images to ensure the diversity among personalized text prompts.
$\epsilon_{\phi}$ HiPer IQ $\downarrow$ IV $\uparrow$ Cosine Sim $\downarrow$ Training time
LoRA No 3.1716 4.2734 0.5383 10.23
LoRA Yes 3.4661 4.7462 0.4726 9.50
Shared tokens Yes 3.5132 4.7181 0.5024 7.16
Table 2: Ablation study of our method.
# HiPer tokens IQ normal-$\downarrow$ IV normal-$\uparrow$ Cosine Sim normal-$\downarrow$
1 4.1406 5.4939 0.5526
5 4.2611 5.4653 0.5580
10 3.7975 5.2350 0.4554
15 4.2976 5.9625 0.3531
20 4.4228 5.8485 0.3983
Table 3: Ablation study on the number of HiPer tokens.
Figure 5: Visual results with different numbers of HiPer tokens.
@@ 5.2 Ablation Study
Number of HiPer tokens $L_{2}$. We vary the number of the HiPer tokens from 1, 5, 10, 15, and 20 using the prompt “Michelangelo style statue of a dog reading news on a cellphone”, Fig. 5 and Tab. 3 show that while altering token length has a subtle impact on the 3D model, the text-to-image model produces images more closely resembling the reference in 2D.
@@
Comparison to LLM-based text prompt augmentation. In addition to personalized text prompts obtained by image-to-text inversion using HiPer [15] as presented in the main paper, here we provide a comparison to a different text prompt sampling technique where the augmented text prompts are generated using large language models (LLMs) from an original text prompt. Particularly, given a text prompt $y$, we use ChatGPT [32] to generate $K$ prompts from $y$ with more details added to describe the object in $y$. For example, with an original text prompt of “A high-quality photo of an ice cream sundae”, we can obtain an augmented text prompt of “A high-quality photo of an ice cream sundae with fresh berries and mint leaves”. We then follow our method in using the $i$-th prompts to condition the $i$-th particle in the VSD loss, respectively. The results are shown in Fig. 6. Compared to our method, LLM-based augmented text prompts also lead to diverse 3D generations, but their quality and diversity remain inferior compared to our image-to-text inversion.
5.3 User Study
We conducted a user study to evaluate the fidelity and diversity of the results obtained by our method and ProlificDreamer [47]. Our survey displays 23 prompts with results of models A and B, where participants are asked to select a model (A or B) that has better quality and diversity in terms of object shape and appearance. Models A and B are shuffled in each question to prevent users from biasing toward a single selection. For each prompt, four video results for each model are displayed in each row. We distributed the survey to users from various backgrounds, including university students, researchers, and general PC/laptop/mobile phone users. In total, we received responses from $52$ participants. After checking all responses and filtering out incomplete/invalid responses, there remains a total of $45$ participants with $1035$ responses. The results of the survey in Tab. 4 show that our method is better in terms of fidelity and diversity than VSD [47], with more than $60\%$ of the responses favoring our results.
Quality Diversity
VSD [47] 39.61% 39.90%
TSD (ours) 60.39% 60.10%
Table 4: User study on the fidelity and diversity of our method compared to ProlificDreamer [47].
Figure 7: Failure cases due to the quality of reference images and capacity of HiPer tokens.
6 Discussion
Limitations. First, despite our model’s ability to enhance the diversity of the existing VSD framework, it still has some limitations. Our models heavily depend on the capacity of the HiPer inversion model [15], and if the reference image is an outlier of what the pretrained diffusion model can generate, HiPer may struggle to accurately invert it, resulting in 3D models with weird shape and appearance. Some failure cases are shown in Fig. 7. Second, as our method is based on VSD, we share its limitations such as the Janus problem. Orthogonal methods that mitigate the Janus problem for VSD should be applicable to our method as well.
@@ 6 Discussion
Conclusion. In this paper, we have successfully introduced a new text-to-3D synthesis method that focuses on diversifying the 3D generation by using 2D reference images and textual inversion to build augmented text prompts for conditioning the optimization. In future work, we plan to experiment with our augmented text embedding technique for other text-to-3D methods [22, 8], and more 3D representations [21].

### v2

@@ 5.1 Comparison with Prior Methods
Qualitative comparison. The comparative results in Fig. 5 demonstrate that our method offers more diversity among particles compared to VSD and SDS. For example, when given “A high-quality ice cream sundae” prompts, VSD tends to collapse into cone-shaped ice cream, while our method is capable of generating glass shapes and other variants. Our 3D models inherit texture and structure from reference images (see Fig. 1), showcasing the potential of transferring 2D diversity to 3D through text prompt augmentation. Perfect inversion by HiPer is not necessary; capturing the essence of reference images suffices for diversity among personalized text prompts. Fig. 6 demonstrates the diversity of our results when applying HiPer to the SDS loss.
@@ 5.2 Ablation Study
In this section, we undertake an ablation study to examine the factors that influence our methods.
Number of HiPer tokens $L_{2}$. We vary the number of the HiPer tokens from 1, 5, 10, 15, and 20 using the prompt “A DSLR photo of a frog wearing a sweater”, Fig. 7 and Tab. 2 show that while altering token length has a subtle impact on the 3D model, the text-to-image model produces images more closely resembling the reference in 2D.
@@
LLM-based text prompt augmentation. In addition to personalized text prompts obtained through image-to-text inversion using HiPer [14], we compare our approach to a different text prompt sampling technique using large language models (LLMs). Given an original text prompt $y$, we employ ChatGPT [37] to generate $K$ prompts from $y$, enriching the description of the object. For instance, starting with “A high-quality photo of an ice cream sundae” as the original prompt, we obtain an augmented prompt like “A high-quality photo of an ice cream sundae with fresh berries and mint leaves”. Subsequently, we utilize each of these prompts to condition the corresponding particle in the VSD loss. The results, shown in Fig. 8, indicate that while LLM-based augmented text prompts also lead to diverse 3D generations, their quality and diversity are inferior compared to our image-to-text inversion method.
5.3 Extension to 3DGS
To expedite training, our approach can be extended to 3D Gaussian Splatting (3DGS) [25]. Since the 3DGS method requires a point cloud to start the optimization, we utilize recent generative text-to-3D approaches such as 3DTopia [16] and Shape-E [22] to obtain the initial shapes for our training. Specifically, we conducted experiments on 3DGS using the same initial shape from 3DTopia and multiple initial shapes from Shape-E. As shown in Figure 9, our method can generate high-quality and diverse 3D renderings using the 3DGS backbone. The 3DGS representation, compared to Instant-NGP, significantly reduces the training time from approximately 7 hours to about 2 hours when using 6 particles.
Figure 9: Our method also achieve remarkable diversity on 3DGS. We also show the initialized point-cloud at the top right corner of each sample.
6 Discussion and Conclusion
Limitations: Despite enhancing the diversity of the existing VSD framework, our model has limitations. It heavily relies on the HiPer inversion, which may struggle with outlier reference images, resulting in 3D models with unusual shapes and appearances. Also, our method shares VSD’s limitations, such as the Janus problem, which can be addressed by orthogonal methods applicable to VSD.
@@ 6 Discussion and Conclusion
Conclusion: In this paper, we have successfully introduced a new text-to-3D synthesis method that focuses on diversifying the 3D generation by using 2D reference images and textual inversion to build augmented text prompts for conditioning the optimization. In future work, we plan to experiment with our augmented text embedding technique for other text-to-3D methods [29, 8], and more 3D representations [25].
Acknowledgment. Binh-Son Hua is supported by Science Foundation Ireland under the SFI Frontiers for the Future Programme (22/FFP-P/11522).